how do i find the system of equations to this problem: julia went to the movies and bought one jumbo popcorn and two chocolate chip cookies for $5. marvin went to the same movie and bought one jumbo popcorn and four chocolate chip cookies for $6.How much does on chocolate chip cookie cost?
PLEASE HELP!!!!

Answer by mananth(16946)

Jumbo Popcorn x
choclate chips y

x+2y=5 .............1
x+4y=6 .............2

multiply (1)by -2
Multiply (2) by 1
-2x-4y=-10
1x+4y= 6
Add the two equations
-x=-4
x=4 Jumbo Popcorn
plug value of x in (1)
4+2y=5
4+2y= 5
y=5-4
y=0.5 choclate chips